# 3Research Center for Information Technology Innovation (CITI), Academia Sinica (AS), Taipei 11529, Taiwan

### v2

@@ Abstract
In this paper, we present an agentic double deep Q-network (DDQN) scheduler for licensed/unlicensed band allocation in New Radio (NR) sidelink (SL) networks. Beyond conventional reward-seeking reinforcement learning (RL), the agent perceives and reasons over a multi-dimensional context that jointly captures queueing delay, link quality, coexistence intensity, and switching stability. A capacity-aware, quality of service (QoS)-constrained reward aligns the agent with goal-oriented scheduling rather than static thresholding. Under constrained licensed bandwidth, the proposed design reduces blocking by up to 87.5% versus threshold policies while preserving throughput, highlighting the value of context-driven decisions in coexistence-limited NR SL systems.
Index Terms:
Sidelink (SL), device-to-device (D2D), unlicensed band, resource allocation, DDQN, agentic AI.
@@ I Introduction
Device-to-device (D2D) communications were introduced in 3GPP LTE Release 12 [1] to support proximity-based services such as vehicle-to-vehicle (V2V) communication. This concept was extended in 5G New Radio (NR) Release 17 [2], where NR sidelink (SL) communications were standardized for cellular vehicle-to-everything (C-V2X) and other emerging applications. In NR SL, resource allocation operates in two modes: a controlled mode, where the base station (gNB) centrally manages allocation, and an autonomous mode, where transmission opportunities are selected by user equipment (UEs) themselves [3]. Detailed protocol procedures for both modes are specified in 3GPP TR 38.331 [4].
@@
These dynamics make resource allocation for NR SL uniquely challenging. Unlike traditional cellular scheduling, which can control all channels, SL-U must share the spectrum with uncontrollable Wi-Fi nodes. Prior approaches have attempted threshold-based scheduling [13, 14], rule-driven heuristics, or Markov-based models [15]. While valuable, these methods lack adaptability to real-time variations in queueing load, channel fading, and coexistence intensity. Moreover, frequent switching between licensed and unlicensed modes can destabilize performance unless carefully constrained, an aspect often overlooked in prior works.
Reinforcement learning (RL) [16] offers a natural solution by enabling agents to adaptively learn from interaction with stochastic environments. Deep Q-networks (DQN) [17, 18] have been widely applied to wireless scheduling, but are prone to Q-value overestimation. Double DQN (DDQN) [19] addresses this limitation by decoupling action selection from evaluation, achieving more stable convergence. Recent advances in agentic AI highlight that agents differ from conventional RL by extending beyond scalar reward optimization toward autonomous, context-driven, and goal-aligned reasoning [20]. Following this perspective, our DDQN-based scheduler embodies agentic decision-making by perceiving heterogeneous context, reasoning across layers, and adapting strategies to maintain QoS under constrained spectrum resources.
Motivated by these trends, we develop an agentic DDQN framework for NR SL that jointly allocates licensed and unlicensed spectrum. Our design integrates queueing delay, channel quality, Wi-Fi coexistence dynamics, and link-switching stability into the state and reward formulation, departing from static heuristics or binary-reward DRL baselines [21, 22]. Unlike conventional RL schedulers that typically optimize a single reward with simplified states (e.g., queue length or binary success), our agent leverages multi-dimensional perception with a capacity-aware, QoS-constrained objective. This enables context-driven, goal-oriented reasoning across licensed/unlicensed modes and avoids pathological behaviors such as overusing a single band or oscillating between modes.
As demonstrated in our experiments, the proposed scheduler achieves up to $87.5\%$ and $82.5\%$ reduction in blocking relative to threshold and random policies, respectively, under constrained licensed bandwidth, while preserving throughput. These results highlight the potential of agentic DDQN-driven learning, consistent with the RL-oriented definition in [20], to deliver QoS-aware, adaptive, and stable scheduling in coexistence-limited NR SL systems.
@@ II Related Work
RL has become a prominent tool for dynamic spectrum access and adaptive scheduling in unlicensed networks [18]. Several studies have applied DQN to address challenges related to NR-U and Wi-Fi coexistence. For example, the authors in [23] proposed an asynchronous multi-channel DQN-based band allocation scheme for NR-U and WiGig systems, focusing on fairness and collision mitigation. The authors in [24] developed a DRL-based framework for joint codebook selection and UE scheduling in mmWave NR-U networks, although their approach did not address mode selection or queueing dynamics.
@@
  • Agentic DDQN scheduler: A DDQN agent that perceives multi-dimensional context (queueing, residual spectrum, channel quality, coexistence, switching stability) and reasons toward QoS goals, departing from conventional reward-only RL.
@@ III System Model
We consider a heterogeneous wireless network in which a 5G NR-U gNB and multiple Wi-Fi access points (APs) coexist in the same region, as shown in Fig. 1. The gNB serves both CC and SL users through licensed and unlicensed spectrum resources, while Wi-Fi nodes independently operate on the 5 GHz unlicensed band.
@@
For SL mode selection, the gNB assigns an SL transmitter-receiver pair and allocates frequency resources. SL-U transmissions compete with Wi-Fi using LBT aligned with IEEE 802.11 CSMA/CA. Channel contention may cause deferrals or blocking, depending on Wi-Fi traffic patterns. Each transmission mode occupies a predefined bandwidth and experiences link-specific path loss and fading. mmWave links (26/28 GHz) suffer from high attenuation and blockage, while the 5 GHz unlicensed band is more susceptible to interference from nearby Wi-Fi devices. The system aims to intelligently assign packets to mode-band pairs that minimize blocking probability while ensuring fairness and efficiency across bands.
@@ VII Conclusion
In this paper, we propose an agentic DDQN-based scheduler for licensed and unlicensed band allocation in NR sidelink networks. Different from vanilla DDQN applications, the design couples rich perception, a goal-aligned reward, and a unified action space, enabling context-driven decisions. Simulations showed up to $87.5\%$ and $82.5\%$ blocking reduction versus threshold-based and random policies at $500$ Mbps licensed bandwidth, while maintaining throughput. These results highlight the promise of agentic AI for robust coexistence, with future work extending to traffic prediction and adaptive LBT tuning.

### v3

# Po-Heng Chou1,3, Pin-Qi Fu2, Walid Saad4, and Li-Chun Wang1,2

@@ Abstract
In this paper, we present an agentic double deep Q-network (DDQN) scheduler for licensed/unlicensed band allocation in New Radio (NR) sidelink (SL) networks. Beyond conventional reward-seeking reinforcement learning (RL), the agent perceives and reasons over a multi-dimensional context that jointly captures queueing delay, link quality, coexistence intensity, and switching stability. A capacity-aware, quality of service (QoS)-constrained reward aligns the agent with goal-oriented scheduling rather than static thresholding. Under constrained bandwidth, the proposed design reduces blocking by up to 87.5% versus threshold policies while preserving throughput, highlighting the value of context-driven decisions in coexistence-limited NR SL networks. The proposed scheduler is an embodied agent (E-agent) tailored for task-specific, resource-efficient operation at the network edge.
@@ I Introduction
Device-to-device (D2D) communications were introduced in 3GPP LTE Release 12 [1] to support proximity-based services such as vehicle-to-vehicle (V2V) communication. This concept was extended in 5G New Radio (NR) Release 17 [2], where NR sidelink (SL) networks were standardized for cellular vehicle-to-everything (C-V2X) and other emerging applications. In NR SL networks, resource allocation operates in two modes: a controlled mode, where the base station (gNB) centrally manages allocation, and an autonomous mode, where transmission opportunities are selected by user equipment (UEs) themselves [3]. Detailed protocol procedures for both modes are specified in 3GPP TR 38.331 [4].
@@
These dynamics make resource allocation for NR SL networks uniquely challenging. Unlike traditional cellular scheduling, which can control all channels, SL-U must share the spectrum with uncontrollable Wi-Fi nodes. Prior approaches have attempted threshold-based scheduling [13, 14], rule-driven heuristics, or Markov-based models [15]. While valuable, these methods lack adaptability to real-time variations in queueing load, channel fading, and coexistence intensity. Moreover, frequent switching between licensed and unlicensed modes can destabilize performance unless carefully constrained, an aspect often overlooked in prior works.
Reinforcement learning (RL) [16] offers a natural solution by enabling agents to adaptively learn from interaction with stochastic environments. Deep Q-networks (DQN) [17, 18] have been widely applied to wireless scheduling, but are prone to Q-value overestimation. Double DQN (DDQN) [19] addresses this limitation by decoupling action selection from evaluation, achieving more stable convergence. Recent advances in agentic AI highlight that agents differ from conventional RL by extending beyond scalar reward optimization toward autonomous, context-driven, and goal-aligned reasoning [20]. Following this perspective, our DDQN-based scheduler constitutes an embodied agent (E-agent) in [20], by perceiving heterogeneous context, reasoning across layers, and adapting strategies to maintain QoS under constrained spectrum resources.
Motivated by these trends, we develop an agentic DDQN framework that instantiates an E-agent for NR SL networks, jointly allocating licensed and unlicensed spectrum. Our design integrates queueing delay, channel quality, Wi-Fi coexistence dynamics, and link-switching stability into the state and reward formulation, departing from static heuristics or binary-reward DRL baselines [21, 22]. Unlike conventional RL schedulers that typically optimize a single reward with simplified states (e.g., queue length or binary success), our agent leverages multi-dimensional perception with a capacity-aware, QoS-constrained objective. This enables context-driven, goal-oriented reasoning across licensed/unlicensed modes and avoids pathological behaviors such as overusing a single band or oscillating between modes.
As demonstrated in our experiments, the proposed scheduler achieves up to $87.5\%$ and $82.5\%$ reduction in blocking relative to threshold and random policies, respectively, under constrained bandwidth, while preserving throughput. These results highlight the potential of agentic DDQN-driven learning, consistent with the RL-oriented definition in [20], to deliver QoS-aware, adaptive, and stable scheduling in coexistence-limited NR SL networks.
@@ II Related Work
RL has become a prominent tool for dynamic spectrum access and adaptive scheduling in unlicensed networks [18]. Several studies have applied DQN to address challenges related to NR-U and Wi-Fi coexistence. For example, the authors in [23] proposed an asynchronous multi-channel DQN-based band allocation scheme for NR-U and WiGig systems, focusing on fairness and collision mitigation. The authors in [24] developed a DRL-based framework for joint codebook selection and UE scheduling in mmWave NR-U networks, although their approach did not address mode selection or queueing dynamics. Compared with foundation-model-based agents (F-agents) and hybrid agents (H-agents) discussed in [20], our design falls into the E-agent category, emphasizing lightweight, domain-specific adaptation for real-time scheduling.
@@
  • Agentic DDQN scheduler as an E-agent: A DDQN agent that perceives multi-dimensional context (queueing, residual spectrum, channel quality, coexistence, switching stability) and reasons toward QoS goals, departing from conventional reward-only RL.
@@ III System Model
We consider a heterogeneous NR SL network where a gNB serves CC and SL users over licensed 26/28 GHz and unlicensed 5 GHz bands, coexisting with Wi-Fi nodes that access the same unlicensed channel via LBT/CSMA-CA, as shown in Fig. 1. Five 3GPP-compliant transmission options are available: CC and SL-L at 26/28 GHz (FR2) and SL-U at 5 GHz under LBT.
@@
Each transmission mode occupies a predefined bandwidth and experiences link-specific path loss and fading. mmWave links (26/28 GHz) suffer from high attenuation and blockage, while the 5 GHz unlicensed band is more susceptible to interference from nearby Wi-Fi devices.
@@ VII Conclusion
In this paper, we propose an agentic DDQN-based scheduler for licensed and unlicensed band allocation in NR sidelink networks. Different from vanilla DDQN applications, the design couples rich perception, a goal-aligned reward, and a unified action space, enabling context-driven decisions. Simulations showed up to $87.5\%$ and $82.5\%$ blocking reduction versus threshold-based and random policies at $500$ Mbps licensed bandwidth, while maintaining throughput. These results highlight the promise of agentic DDQN as an E-agent for QoS-aware, adaptive, and stable scheduling in coexistence-limited NR SL networks, with future extensions toward H-agents that fuse embodied learning with foundation-model reasoning.